# Understanding Adam Requires Better Rotation Dependent Assumptions

### v2

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities as their scale grows [Brown et al., 2020, Kaplan et al., 2020]. However, this unprecedented growth in model scale has led to a proportional increase in the economic [Dong and Xie, 2024, Sharir et al., 2020, Varoquaux et al., 2024] and environmental [Luccioni et al., 2023, 2019] costs associated with their training. Despite this clear motivation, Adaptive Moment Estimation (Adam) [Kingma and Ba, 2015] has persisted as the go-to optimizer especially for language models, with only minor modifications such as AdamW [Loshchilov and Hutter, 2019] becoming widely adopted since Adam’s inception. This success has prompted extensive research to provide theoretical justification for Adam’s performance. While the original convergence proof for Adam was later found to be flawed [Rubio, 2017], recent studies have proposed rigorous convergence proofs under plausible assumptions [Li et al., 2023b, Chen et al., 2019, Défossez et al., 2022].
However, these proofs do not elucidate Adam’s advantages over SGD when training transformer models [Vaswani et al., 2017]. Numerous works attempted to explain Adam’s superiority, employing diverse assumptions and analytical frameworks [Zhou et al., 2024, Pan and Li, 2022, Zhang et al., 2024, Kunstner et al., 2024]. The heterogeneity of these approaches has led to a lack of consensus on which theoretical explanations most accurately capture the fundamental mechanisms underlying Adam’s improved performance. For instance, Zhang et al. [2020] suggests it stems from enhanced robustness to heavy-tailed noise, while  Kunstner et al. [2023] argues it plays no role.
@@
This study focuses on a fundamental distinction between Adam and SGD: Adam’s dependency on the coordinate system. SGD is rotation-equivariant, meaning if the loss landscape is rotated, the resulting optimization trajectories from SGD will be the same up to that rotation. In contrast, Adam produces substantially different trajectories. Although the sensitivity of adaptive methods to rotations of the parameter space is well established [Duchi et al., 2011], it remains unclear what properties of a rotation benefit or hinder performance, particularly in practical neural network training.
@@
Beyond theoretical motivations, recent studies have shown that applying rotations to the canonical basis can significantly enhance the performance of Adam and other optimizers [Vyas et al., 2025, Gupta et al., 2018, Jordan et al., 2024]. However, these rotations are often designed based on intuition and heuristics rather than systematically understanding their impact. By identifying the key properties that make a basis advantageous for Adam, we can develop more principled approaches to constructing optimal rotations that may outperform the existing rotation-based methods.
@@
Section˜4 examines three rotation-dependent assumptions used in literature in this context: $L_{\infty}$ bounded gradients, Hessian block-diagonality, and $L_{\infty}$-smoothness. Our analysis reveals that none of these conditions fully capture Adam’s behaviour under rotations. Recently, Muon [Jordan et al., 2024] achieved strong performance by approximating orthogonalized gradients for each layer. Inspired by this, we measure the orthogonality of Adam’s weight matrix updates for each layer (up to scalar multiplication) and find it closely aligns with performance across various rotations.
@@
    Analysis of Adam’s sensitivity to various scopes of parameter space rotations in neural network training: We conduct in Section˜3.1 an empirical study demonstrating Adam’s sensitivity to random parameter space rotations in practical training. We found a clear correlation between rotation scope (e.g., global, layer-wise) and performance, where a broader scope leads to greater degradation. In Section˜3.2 we also employ a structured, SVD-based rotation inspired by Zhao et al. [2024] that improves Adam’s convergence.
@@
    Verifying a better rotation-dependent quantity: Given that SVD-based rotations improve performance, we analyze the singular values of the layer updates and find that their orthogonality is a strong predictor of Adam’s performance across different bases. We also draw a connection to the Muon optimizer [Jordan et al., 2024], which approximates orthogonalized updates, and provide additional empirical support for its underlying intuition.
@@ 2.2 Training Neural Networks in Rotated Parameter Spaces
Numerical considerations. Neural network training is sensitive to numerical precision [Li et al., 2018, Wang et al., 2018, Sun et al., 2022], and it is crucial to ensure that rounding errors from rotations do not significantly confound the impact of the rotation. In particular, we apply rotations in single precision, and we refrain from using FlashAttention [Dao et al., 2022], which was found to increase numeric deviations [Golden et al., 2024]. We validate our methodology with ablations on SGD rotation equivalence, various rotation dimensions, and the use of FlashAttention in Appendix˜A.
@@ 3.1 Random Rotations
  • Language modeling (GPT-2, Fig. 1a): 124M-parameter decoder-only Transformer [Radford et al., 2019] trained on OpenWebText [Gokaslan and Cohen, 2019].
  • Image classification (ViT, Fig. 1b): 22M-parameter Vision Transformer (ViT/S) [Dosovitskiy et al., 2021] evaluated on ImageNet-1K [Deng et al., 2009].
  • Image classification (ResNet, Fig. B): ResNet-50 [He et al., 2016] on ImageNet-1K, where SGD often outperforms Adam [Keskar and Socher, 2017, Wilson et al., 2017].
@@
    Output-wise rotations show no degradation across all settings, with GPT2 even slightly improving. This suggests that Adam’s adaptivity within output neurons is minimal, supporting recent approaches to reduce redundancy in Adam’s second moments [Zhang et al., 2025].
Previous works [Zhang et al., 2024, 2025] have highlighted the heterogeneity across different parameter block types in Transformer architectures. In Appendix˜C, we restrict rotations to specific parameter types and study their individual impact on Adam’s rotation sensitivity.
@@ 3.2 Investigating the Performance-Improving Rotations
Inspired by GaLore [Zhao et al., 2024], which uses low-rank Singular Value Decomposition (SVD) to compress optimizer states, we extend this concept to full-rank SVD to rotate the parameter space. Our approach decomposes the gradient matrix $\mathbf{G}$ of each layer into $\mathbf{G}=\mathbf{U}\mathbf{S}\mathbf{V}^{\top}$. This decomposition yields a natural rotation of the parameter space through the transformation $\mathbf{G}\rightarrow\mathbf{U}^{\top}\mathbf{G}\mathbf{V}$, which corresponds to an output-wise rotation (via $\mathbf{U}^{\top}$) and an input-wise rotation (via $\mathbf{V}$). We train a GPT2 model under the same conditions as in Section˜3.1, but in this SVD-rotated space. We update the SVD decompositions every 250 steps.
@@ 4.1 Adequacy of existing assumptions in theoretical frameworks
Reddi et al. [2018], Kingma and Ba [2015] assume a bound on the $L_{\infty}$ norm of stochastic gradients,
@@
$L_{\infty}$-smoothness was recently shown to guarantee the convergence of Adam, and presented as a potential key property of the basis [Xie et al., 2025, Balles et al., 2020]. We first remember its definition.
@@
Given the challenges in directly estimating the $L_{\infty}$-smoothness constant, Xie et al. [2025] proposed using the (1,1)-norm of the Hessian as a surrogate, defined as:
@@
Under global rotations, we confirm the $(1,1)$-norm degradation reported in [Xie et al., 2025], while SVD rotations improve it in line with Adam’s performance gains, suggesting a link between this geometric property and optimizer efficiency. Output-wise rotations yield slight performance gains but reduced $(1,1)$-norm, indicating that the metric does not capture all relevant factors. Overall, the $(1,1)$-norm shows promise as a rotation-sensitive indicator, especially under global and SVD rotations, but its limitations motivate the development of refined or complementary measures.
@@
A common hypothesis for understanding Adam’s behaviour is that the Hessian is well-approximated by a block-diagonal matrix [Zhang et al., 2024]. Then, random rotations likely disrupt block-diagonality and hinder convergence, while rotations within diagonal blocks preserve the structure, explaining the stable performance under output-wise rotations.
@@
where $\mathbf{e_{i}}$ denotes the $i^{\textit{th}}$ canonical basis vector. The vector $\mathbf{r_{i}}$ represents the average of the $i$-th row of the stochastic Hessian over $k$ minibatches. As $k$ increases, $\mathbf{r_{i}}$ converges to the true Hessian row. We set $k=5000$ in the experiments, and use efficient Hessian-vector products  [Dagréou et al., 2024].
@@
Given this approximately block-diagonal Hessian structure, previous work [Zhang et al., 2024] proposes a strict block-diagonal approximation, assuming that the off-diagonal elements are negligible. We further investigate whether this simplification can accurately reflect the local geometry by assessing the practical implications of the block diagonal Hessian structure. We evaluate how each block contributes to gradient variations for a given small direction $\delta w$ via:
@@ 4.2 Orthogonality of layer updates up to scalar factor
In Section˜3.2 we find that SVD-based rotations improve Adam’s performance. We also discuss in Appendix˜E connections with the recently proposed Muon optimizer [Jordan et al., 2024], which achieves strong performance by performing updates in the orthogonalized first moment direction.
Scaled (semi-)orthogonality. Since orthogonality is defined only for square matrices, we adopt a relaxed notion for rectangular matrices. Specifically, we say that a rectangular matrix is orthogonal if all its singular values are either 1 or 0 (this notion is commonly referred to as semi-orthogonality [Abadir and Magnus, 2005]). Moreover, we say that a matrix is a scaled orthogonal matrix if its eigenvalues are either $\alpha$ or 0, where $\alpha$ is the scaling parameter. To measure the scaled orthogonality of a matrix $\mathbf{A}$, we will use the coefficient of variation of its singular values $s_{i}$,
@@
In Figure 10, we show the CV across time for each layer type, averaged over the depth. See Section˜C.4 for full results per layer and over time. Notably, CV drops right after the SVD rotation is recomputed every 250 steps. This offers insight into the frequency tradeoff of SVD rotations, where more frequent updates improve performance but introduce computational overhead. Overall, our analysis suggests that update orthogonality strongly correlates with optimizer performance, supporting the approaches of Muon [Jordan et al., 2024] and SOAP [Vyas et al., 2025] and opening new avenues for rotation-aware theoretical frameworks.
While a robust theoretical analysis of Muon’s update rule remains open, attempts have been made to characterize in what context an orthogonal update is optimal. Bernstein [2025] argues that for linear layers, the orthogonalized update controls the scale at which the weight matrices can scale features. They argue this encourages stable optimization and can limit the need for normalization layers. As we discuss in Appendix˜E, after simplification and SVD rotation the Adam update simplifies to dividing the singular values by their magnitude. When rotated back, this results in the same update as Muon through a different mechanism, which is consistent with Muon’s often superior performance.
@@ 5 Related Work
Shampoo [Gupta et al., 2018] first demonstrated the benefits of optimizing under rotations by running AdaGrad under regularly updated rotations, identifying the singular values of the gradient matrices. More recently, SOAP [Vyas et al., 2025] improved Shampoo by applying it to Adam and appropriately updating the moments at each change of basis. Muon [Jordan et al., 2024] concurrently explores a similar approach, but removes the need to explicitly store rotation matrices and instead orthogonalizes gradient matrices with a fast matrix iteration. While these methods show promising empirical improvements via heuristics, our work highlights the importance of developing better theoretical tools to understand their success.
On the theoretical side, we consider [Xie et al., 2025] to be the closest related study, showing that Adam converges more slowly with a randomly rotated loss landscape. They provide convergence analysis based on $L_{\infty}$ geometry, demonstrating that this yields a better empirical smoothness constant for GPT-2 models. While their work offers valuable theoretical insights, our study takes a more experimental stance. We aim to paint a comprehensive picture of Adam’s behaviour under a spectrum of rotations, from random to structured transformations, and evaluate how existing rotation-invariant assumptions correlate with Adam’s performance. Notably, Balles et al. [2020] also provides relevant insights through the lens of sign gradient descent.
@@
Our work casts light on the critical interactions between Adam and the coordinate system, contributing to a growing body of research on Adam’s behaviour and convergence. Recent works have attributed Adam’s success to the heterogeneous block-diagonal structure of its Hessian [Zhang et al., 2024], though we find this assumption to be unrealistic. Others have improved convergence guarantees: Défossez et al. [2022] and Guo et al. [2022] offered simplified and novel derivations, Zhang et al. [2022] argued that vanilla Adam converges without modification, Zhou et al. [2024] provided a general convergence analysis for adaptive methods in non-convex settings, and Li et al. [2023b] proposed a convergence proof for Adam without relying on globally bounded gradients. Li et al. [2023a] developed a convergence analysis based on generalized smoothness conditions, and Hübler et al. [2023] proposed parameter-agnostic convergence results under these relaxed conditions. Finally, lower bounds for non-convex optimization were established by Arjevani et al. [2019], with Wang et al. [2023] addressing the gap between upper and lower bounds for Adam’s iteration complexity.
@@
Prior works have attempted to justify Adam’s advantages over SGD. Zhang et al. [2020], Zhou et al. [2024] suggest SGD suffers more from heavy-tailed noise, with Adam converging faster when gradients are sparse. However, Kunstner et al. [2023] found that noise reduction through larger batch sizes benefits Adam but not SGD. Additionally, Kunstner et al. [2024] ties Adam’s advantage in language models to ill-conditioning caused by heavy-tailed class imbalance, and Pan and Li [2022] to directional sharpness.

### v3

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities as their scale grows [NEURIPS2020_1457c0d6, kaplan2020scaling]. However, this unprecedented growth in model scale has led to a proportional increase in the economic [dong2024large, sharir2020cost, varoquaux2024hypesustainabilitypricebiggerisbetter] and environmental [JMLR:v24:23-0069, luccioni2019quantifying] costs associated with their training. Despite this clear motivation, Adaptive Moment Estimation (Adam) [DBLP:journals/corr/KingmaB14] has persisted as the go-to optimizer especially for language models, with only minor modifications such as AdamW [DBLP:conf/iclr/LoshchilovH19] becoming widely adopted since Adam’s inception. This success has prompted extensive research to provide theoretical justification for Adam’s performance. While the original convergence proof for Adam was later found to be flawed [rubio2017convergence], recent studies have proposed rigorous convergence proofs under plausible assumptions [NEURIPS2023_a3cc5012, chen2018on, defossez2022a].
However, these proofs do not elucidate Adam’s advantages over SGD when training transformer models [10.5555/3295222.3295349]. Numerous works attempted to explain Adam’s superiority, employing diverse assumptions and analytical frameworks [zhou2024on, pan2022toward, DBLP:journals/corr/abs-2402-16788, kunstner2024heavytailedclassimbalanceadam]. The heterogeneity of these approaches has led to a lack of consensus on which theoretical explanations most accurately capture the fundamental mechanisms underlying Adam’s improved performance. For instance, NEURIPS2020_b05b57f6 suggests it stems from enhanced robustness to heavy-tailed noise, while  kunstner2023noise argues it plays no role.
@@
This study focuses on a fundamental distinction between Adam and SGD: Adam’s dependency on the coordinate system. SGD is rotation-equivariant, meaning if the loss landscape is rotated, the resulting optimization trajectories from SGD will be the same up to that rotation. In contrast, Adam produces substantially different trajectories. Although the sensitivity of adaptive methods to rotations of the parameter space is well established [duchi11a], it remains unclear what properties of a rotation benefit or hinder performance, particularly in practical neural network training.
@@
Beyond theoretical motivations, recent studies have shown that applying rotations to the canonical basis can significantly enhance the performance of Adam and other optimizers [vyas2025soap, 47079, jordan2024muon]. However, these rotations are often designed based on intuition and heuristics rather than systematically understanding their impact. By identifying the key properties that make a basis advantageous for Adam, we can develop more principled approaches to constructing optimal rotations that may outperform the existing rotation-based methods.
@@
Section˜4 examines three rotation-dependent assumptions used in literature in this context: $L_{\infty}$ bounded gradients, Hessian block-diagonality, and $L_{\infty}$-smoothness. Our analysis reveals that none of these conditions fully capture Adam’s behaviour under rotations. Recently, Muon [jordan2024muon] achieved strong performance by approximating orthogonalized gradients for each layer. Inspired by this, we measure the orthogonality of Adam’s weight matrix updates for each layer (up to scalar multiplication) and find it closely aligns with performance across various rotations.
@@
    Analysis of Adam’s sensitivity to various scopes of parameter space rotations in neural network training: We conduct in Section˜3.1 an empirical study demonstrating Adam’s sensitivity to random parameter space rotations in practical training. We found a clear correlation between rotation scope (e.g., global, layer-wise) and performance, where a broader scope leads to greater degradation. In Section˜3.2 we also employ a structured, SVD-based rotation inspired by zhao2024galore that improves Adam’s convergence.
@@
    Verifying a better rotation-dependent quantity: Given that SVD-based rotations improve performance, we analyze the singular values of the layer updates and find that their orthogonality is a strong predictor of Adam’s performance across different bases. We also draw a connection to the Muon optimizer [jordan2024muon], which approximates orthogonalized updates, and provide additional empirical support for its underlying intuition.
@@ 2.2 Training Neural Networks in Rotated Parameter Spaces
Numerical considerations. Neural network training is sensitive to numerical precision [li2018explorationnumericalprecisiondeep, NEURIPS2018_335d3d1c, sun2022surprising], and it is crucial to ensure that rounding errors from rotations do not significantly confound the impact of the rotation. In particular, we apply rotations in single precision, and we refrain from using FlashAttention [NEURIPS2022_67d57c32], which was found to increase numeric deviations [golden2024flashattentionstable]. We validate our methodology with ablations on SGD rotation equivalence, various rotation dimensions, and the use of FlashAttention in Appendix˜A.
@@ 3.1 Random Rotations
  • Language modeling (GPT-2, Fig. 1a): 124M-parameter decoder-only Transformer [radford2019language] trained on OpenWebText [Gokaslan2019OpenWeb].
  • Image classification (ViT, Fig. 1b): 22M-parameter Vision Transformer (ViT/S) [dosovitskiy2021an] evaluated on ImageNet-1K [deng2009imagenet].
  • Image classification (ResNet, Fig. B): ResNet-50 [7780459] on ImageNet-1K, where SGD often outperforms Adam [keskar2017improvinggeneralizationperformanceswitching, NIPS2017_81b3833e].
@@
    Output-wise rotations show no degradation across all settings, with GPT2 even slightly improving. This suggests that Adam’s adaptivity within output neurons is minimal, supporting recent approaches to reduce redundancy in Adam’s second moments [zhang2024adamminiusefewerlearning].
Previous works [DBLP:journals/corr/abs-2402-16788, zhang2024adamminiusefewerlearning] have highlighted the heterogeneity across different parameter block types in Transformer architectures. In Appendix˜C, we restrict rotations to specific parameter types and study their individual impact on Adam’s rotation sensitivity.
@@ 3.2 Investigating the Performance-Improving Rotations
Inspired by GaLore [zhao2024galore], which uses low-rank Singular Value Decomposition (SVD) to compress optimizer states, we extend this concept to full-rank SVD to rotate the parameter space. Our approach decomposes the gradient matrix $\mathbf{G}$ of each layer into $\mathbf{G}=\mathbf{U}\mathbf{S}\mathbf{V}^{\top}$. This decomposition yields a natural rotation of the parameter space through the transformation $\mathbf{G}\rightarrow\mathbf{U}^{\top}\mathbf{G}\mathbf{V}$, which corresponds to an output-wise rotation (via $\mathbf{U}^{\top}$) and an input-wise rotation (via $\mathbf{V}$). We train a GPT2 model under the same conditions as in Section˜3.1, but in this SVD-rotated space. We update the SVD decompositions every 250 steps.
@@ 4.1 Adequacy of existing assumptions in theoretical frameworks
j.2018on, DBLP:journals/corr/KingmaB14 assume a bound on the $L_{\infty}$ norm of stochastic gradients,
@@
$L_{\infty}$-smoothness was recently shown to guarantee the convergence of Adam, and presented as a potential key property of the basis [xie2025adam, balles2020geometrysigngradientdescent]. We first remember its definition.
@@
Given the challenges in directly estimating the $L_{\infty}$-smoothness constant, xie2025adam proposed using the (1,1)-norm of the Hessian as a surrogate, defined as:
@@
Under global rotations, we confirm the $(1,1)$-norm degradation reported in [xie2025adam], while SVD rotations improve it in line with Adam’s performance gains, suggesting a link between this geometric property and optimizer efficiency. Output-wise rotations yield slight performance gains but reduced $(1,1)$-norm, indicating that the metric does not capture all relevant factors. Overall, the $(1,1)$-norm shows promise as a rotation-sensitive indicator, especially under global and SVD rotations, but its limitations motivate the development of refined or complementary measures.
@@
A common hypothesis for understanding Adam’s behaviour is that the Hessian is well-approximated by a block-diagonal matrix [DBLP:journals/corr/abs-2402-16788]. Then, random rotations likely disrupt block-diagonality and hinder convergence, while rotations within diagonal blocks preserve the structure, explaining the stable performance under output-wise rotations.
@@
where $\mathbf{e_{i}}$ denotes the $i^{\textit{th}}$ canonical basis vector. The vector $\mathbf{r_{i}}$ represents the average of the $i$-th row of the stochastic Hessian over $k$ minibatches. As $k$ increases, $\mathbf{r_{i}}$ converges to the true Hessian row. We set $k=5000$ in the experiments, and use efficient Hessian-vector products  [dagréou2024howtocompute].
@@
Given this approximately block-diagonal Hessian structure, previous work [DBLP:journals/corr/abs-2402-16788] proposes a strict block-diagonal approximation, assuming that the off-diagonal elements are negligible. We further investigate whether this simplification can accurately reflect the local geometry by assessing the practical implications of the block diagonal Hessian structure. We evaluate how each block contributes to gradient variations for a given small direction $\delta w$ via:
@@ 4.2 Orthogonality of layer updates up to scalar factor
In Section˜3.2 we find that SVD-based rotations improve Adam’s performance. We also discuss in Appendix˜E connections with the recently proposed Muon optimizer [jordan2024muon], which achieves strong performance by performing updates in the orthogonalized first moment direction.
Scaled (semi-)orthogonality. Since orthogonality is defined only for square matrices, we adopt a relaxed notion for rectangular matrices. Specifically, we say that a rectangular matrix is orthogonal if all its singular values are either 1 or 0 (this notion is commonly referred to as semi-orthogonality [abadir2005matrix]). Moreover, we say that a matrix is a scaled orthogonal matrix if its eigenvalues are either $\alpha$ or 0, where $\alpha$ is the scaling parameter. To measure the scaled orthogonality of a matrix $\mathbf{A}$, we will use the coefficient of variation of its singular values $s_{i}$,
@@
In Figure 10, we show the CV across time for each layer type, averaged over the depth. See Section˜C.4 for full results per layer and over time. Notably, CV drops right after the SVD rotation is recomputed every 250 steps. This offers insight into the frequency tradeoff of SVD rotations, where more frequent updates improve performance but introduce computational overhead. Overall, our analysis suggests that update orthogonality strongly correlates with optimizer performance, supporting the approaches of Muon [jordan2024muon] and SOAP [vyas2025soap] and opening new avenues for rotation-aware theoretical frameworks.
While a robust theoretical analysis of Muon’s update rule remains open, attempts have been made to characterize in what context an orthogonal update is optimal. bernstein2025deriving argues that for linear layers, the orthogonalized update controls the scale at which the weight matrices can scale features. They argue this encourages stable optimization and can limit the need for normalization layers. As we discuss in Appendix˜E, after simplification and SVD rotation the Adam update simplifies to dividing the singular values by their magnitude. When rotated back, this results in the same update as Muon through a different mechanism, which is consistent with Muon’s often superior performance.
@@ 5 Related Work
Shampoo [47079] first demonstrated the benefits of optimizing under rotations by running AdaGrad under regularly updated rotations, identifying the singular values of the gradient matrices. More recently, SOAP [vyas2025soap] improved Shampoo by applying it to Adam and appropriately updating the moments at each change of basis. Muon [jordan2024muon] concurrently explores a similar approach, but removes the need to explicitly store rotation matrices and instead orthogonalizes gradient matrices with a fast matrix iteration. While these methods show promising empirical improvements via heuristics, our work highlights the importance of developing better theoretical tools to understand their success.
On the theoretical side, we consider [xie2025adam] to be the closest related study, showing that Adam converges more slowly with a randomly rotated loss landscape. They provide convergence analysis based on $L_{\infty}$ geometry, demonstrating that this yields a better empirical smoothness constant for GPT-2 models. While their work offers valuable theoretical insights, our study takes a more experimental stance. We aim to paint a comprehensive picture of Adam’s behaviour under a spectrum of rotations, from random to structured transformations, and evaluate how existing rotation-invariant assumptions correlate with Adam’s performance. Notably, balles2020geometrysigngradientdescent also provides relevant insights through the lens of sign gradient descent.
@@
Our work casts light on the critical interactions between Adam and the coordinate system, contributing to a growing body of research on Adam’s behaviour and convergence. Recent works have attributed Adam’s success to the heterogeneous block-diagonal structure of its Hessian [DBLP:journals/corr/abs-2402-16788], though we find this assumption to be unrealistic. Others have improved convergence guarantees: defossez2022a and guo2022novelconvergenceanalysisalgorithms offered simplified and novel derivations, NEURIPS2022_b6260ae5 argued that vanilla Adam converges without modification, zhou2024on provided a general convergence analysis for adaptive methods in non-convex settings, and NEURIPS2023_a3cc5012 proposed a convergence proof for Adam without relying on globally bounded gradients. li2023convex developed a convergence analysis based on generalized smoothness conditions, and hubler2023parameteragnostic proposed parameter-agnostic convergence results under these relaxed conditions. Finally, lower bounds for non-convex optimization were established by arjevani2022lowerboundsnonconvexstochastic, with wang2023closing addressing the gap between upper and lower bounds for Adam’s iteration complexity.
@@
Prior works have attempted to justify Adam’s advantages over SGD. NEURIPS2020_b05b57f6, zhou2024on suggest SGD suffers more from heavy-tailed noise, with Adam converging faster when gradients are sparse. However, kunstner2023noise found that noise reduction through larger batch sizes benefits Adam but not SGD. Additionally, kunstner2024heavytailedclassimbalanceadam ties Adam’s advantage in language models to ill-conditioning caused by heavy-tailed class imbalance, and pan2022toward to directional sharpness.